One phone company charges $0.29 per minute for long-distance phone calls. A second company charges only $0.26 per minute for long-distance calls, but adds a monthly fee of $2.97. Find the number of minutes of long-distance calling for which the total bill from either company would be the same. Please round your answers to the nearest hundredth if necessary.
Answer by mananth(16949)

Let x be the number of calls
company B 2.97 + 0.26 x
company A 0 + 0.29 x
company B = company A
0 + 0.29 x = 2.97 + 0.26 x
0 + 0.29 x = 2.97 + 0.26 x
0 - 2.97 = 0.26 x -0.29 x
-2.97 < -0.03 x
99 = x
x should be 99 calls